A battery consisting of three identical cells, each with internal resistance of 1.0 Ω is connected in a series with


5.0 Ω and 3.0 Ω resistors. The voltage across the 5 Ω resistor is 5V. Determine the (a) current flowing through the circuit, (b) voltage across the 3 Ω resistor, (C) terminal voltage of the cell, and (d) electromotive force of the cell.

1
Expert's answer
2022-03-11T10:47:25-0500

Solution.

"r_1=1.0 Om;"

"R_1=5.0 Om;"

"V_1=5.0V;"

"R_2=3.0Om;"

"a) I=\dfrac{V_1}{R_1};"

"I=\dfrac{5.0}{5.0}=1.0A;"

"b)V_2=IR_2;"

"V_2=1.0\sdot3.0=3.0V;"

"c)R=R_1+R_2;"

"R=5.0+3.0=8.0Om;"

"V=IR;"

"V=1.0\sdot8.0=8.0V;"

"d) EMF=I(R+r); r=3r_1;"

"EMF=1.0(8.0+3.0)=11V;"

"EMF_1=\dfrac{EMF}{3};"

"EMF_1=\dfrac{11}{3}=3.7V;"

Answer: "a) I=1.0A;"

"b)V_2=3.0V;"

"c)V=8.0V;"

"d)EMF_1=3.7V."